# What If Moderation Didn’t Mean Suppression? A Case for Personalized Content Transformation

### v2

@@ 6.2.3. Quantitative Alignment with User Preferences
To complement our qualitative findings, we modeled the probability that participants would select the system’s highest-ranked intervention over a nearby alternative—the randomly selected intervention from ranks two through five. We use an intercept-only model to predict our dependent variable, System Choice Favored, which is a binary variable equal to 1 if the participant selected the system’s choice, and 0 otherwise. We use the glmer function from the lme4 package(lme4) in R. For image interventions, the estimated intercept was positive and statistically significant (log-odds = 0.53, $p=.033$), corresponding to a 63% probability that participants chose the system’s top candidate over the close contenders. For text interventions, the alignment was stronger (log-odds = $1.33$, p<.001{p}<.001), corresponding to a 79.1% probability.
@@ 7.1. Transformation Enables Access, Not Avoidance
Empowering users to take moderation decisions into their own hands might raise concerns about whether this agency can trap them in filter bubbles(pariser2011filter). DIY-MOD can support a different dynamic. By transforming how content appears rather than removing it entirely, the system enables users to engage with content they would otherwise avoid completely.
@@
At the same time, this approach surfaces deeper tensions: Who decides what users should or should not engage with, and why does that decision need to be made in a top-down, prescriptive manner? The First Amendment is often framed as a protection of speech. Legal scholars have similarly argued for the right to freedom of listening and the right against compelled listening (corbin2009first). While these protections do not extend to platforms as private entities, the principles that they represent remain relevant. Our studies show that content that appears benign to most can be experienced as harmful by others. This variability complicates any universal judgment about what content is “safe” or “appropriate”—at least outside the categories codified by law. And it raises a critical question: Who gets to decide which experiences of harm are valid and which users are entitled to protection?
@@ 7.2. The Case for Platform-Integrated Personalization
While DIY-MOD demonstrates personalized moderation is viable as middleware (fukuyama2021), its architecture requires intercepting, externally processing, and then scoring generated transformations. These steps introduce unavoidable latency and complicate implementation.
Platforms however, could avoid these bottlenecks entirely. With direct content access, computational resources, and existing user behavioral signals, they could implement personalized moderation far more efficiently. Platforms already use this infrastructure for ad targeting and safety interventions like suicide prevention (facebook_suicide_1; facebook_suicide_2). The same systems could also power consensual, user-directed personalized content transformation.
@@
Our approach gives rise to a natural question: Which kinds of content should be addressed through platform-led moderation, and which might be better suited to user-driven transformation? Platforms are widely expected to bear responsibility for addressing harms that have broad, collective consequences such as coordinated manipulation campaigns or abuse like hate speech or targeted harassment that fosters hostile environments. Yet even here, the boundaries are far from settled. The same activity can be cast as harassment or as activism, as disinformation or as political dissent. Platforms’ judgments about what to enforce are never neutral. They are shaped by political pressures, business incentives, and cultural norms (diaz2021double; jahanbakhsh2023thesis).
@@ 7.4. Application to Civic Discourse
This transformation-based approach may also hold promise for other types of content not brought up by our study participants. One can imagine its application to civic or ideological discourse: when valuable perspectives are obscured by hostile or inflammatory framing, repackaging that content (e.g., by softening antagonistic language, reducing confrontational cues, or highlighting shared values), could make it easier to engage with. It is also worth noting that there is no clear consensus that exposure to counter-attitudinal information is always beneficial. In fact, a body of research shows that such exposure, especially when unfiltered or confrontational, can exacerbate polarization and deepen animosity (tornberg2022digital_echo; bail2022breaking). This calls into question that more exposure is always better, and highlights the need for systems that support constructive engagement, rather than unmediated confrontation (bail2022breaking). Altering the tone or phrasing of civic or political content through transformations like ours may offer a path toward openness without overwhelm. It is also worth considering who is most likely to adopt such tools. Prior work suggests challenge-averse individuals may use personalization to reinforce existing views (munson2010presenting). However, those findings reflect systems that offer a binary choice (e.g., show/hide). By contrast, transformation reduces the affective load of counter-attitudinal content. It is plausible that this, unlike binary filtering, might enable engagement from users who would otherwise disengage completely. Investigating these adoption patterns remains a critical area for future research.
Prior work on bridging divides has largely focused on identifying and ranking content that resonates across ideological lines (ovadya2022bridging; stray2021designing). While valuable, such approaches depend on the availability of naturally bridgeable posts. But what if we could make more content bridgeable? This of course, raises important questions about authenticity and message integrity: How much transformation is too much? When does reframing a message become misrepresentation? Future work must address these ethical boundaries.
@@ 7.5. Therapeutic Grounding and Applications of Our Work
DIY-MOD’s design is inspired by exposure therapy: enabling safer engagement with a weakened or modified version of a stimulus, rather than avoidance (rothbaum2002exposure; abramowitz2019exposure). This principle is reflected in our palette of interventions. Our transformations are designed to function as weakened versions of the original content, and the ability to easily adjust sensitivity levels and filter durations echoes the concept of graduated exposure.
While DIY-MOD is not a clinical intervention in its current form, its combination of weakened content, adjustable sensitivities, and time-bounded filters hints at how similar mechanisms could eventually complement therapeutic practices. For example, these knobs could let users (and clinicians) gradually decrease intervention intensity over time, if they so choose, rather than assuming static sensitivities. This direction aligns with calls (wang2025mentally) for practitioner-assisted AI systems that support personalized, adaptive exposure in everyday contexts.
@@ 7.6.1. Architectural Safeguards:
Using AI to interpret personal sensitivities carries inherent risks, as system failures can expose users to content they explicitly sought to avoid. One way we mitigate these risks is by implementing a two-stage cascade architecture: a pruner and then a scorer VLM-judge evaluate content using a consistent rubric. Such multi-stage evaluation strengthens the reliability of LLM/VLM-judge based pipelines (badshah2024reference; gu2024survey). We manually audited sampled content and found that all sampled text edits and approximately 89% of sampled image edits were successfully transformed, with false obfuscations (edits that did not match the viewer’s filter description) in only about 7% of transformed images (Section˜A.5).
@@
Privacy considerations are fundamental when handling data tied to personal sensitivities. DIY-MOD adopts a “privacy-by-design” approach (cavoukian2009privacy) that prioritizes user control and data minimization. It provides full functionality without an account and offers user-controlled portability of filters through an anonymous export/import feature. By default, it uses local-first storage and practices data minimization, sending only public in-feed content for ephemeral analysis. Looking ahead, portions of the analysis can migrate to lightweight on-device models to further reduce reliance on server-side processing.
@@ 8. Limitations and Future Directions
Improving Selection Model. Our findings reveal that the effectiveness of an intervention is context-dependent, and that successful transformations manage a user’s cognitive state by providing closure. Our system’s selection framework, in contrast, relied on a limited user model derived from filter descriptions and lacked an explicit model of cognitive closure or the user’s immediate context. Improving its ability to select transformations that provide such closure is an important area for future work. The principles and preference data from this study provide a direct path to do so by aligning the selector with human preferences using alignment approaches such as DPO (rafailov2023direct).
Long-Term Psychological Implications. Our study focused on users’ immediate experiences, which raises a longitudinal question: do personalized transformations support long-term wellbeing, or can they enable patterns of unhealthy avoidance? There is good reason to shield users from repeated triggers, as re-exposure may exacerbate anxiety (markowitz2020exposure). At the same time, indefinite avoidance may also have adverse psychological and social effects. Future work should explore how to design for this tension perhaps by incorporating features that allow users to optionally and safely decrease intervention intensity over time, aligning the tool more closely with therapeutic goals.
@@
Technical and Platform Dependencies. While DIY-MOD’s core logic is generalizable, the content interception requires platform-specific adapters, and extending DIY-MOD to video-first (e.g., TikTok) platforms would require additional engineering effort. Further, our reliance on commercial VLMs means their capabilities, biases, and safety filters can affect performance. We detail specific failure modes and mitigations in Section˜A.5; developing benchmarks to quantify edit failures and over-detection rates remains a potential future work.
Scalability & Latency. Our middleware approach also introduces computational and monetary costs. Our optimizations (detailed in Appendix˜A) keep the per-batch latency within a usable range for real-time interaction, but high cost associated with LLM/VLM usage remains a scalability constraint. These constraints highlight the efficiency gains of a platform-integrated approach (Section˜7.2). Future work could also explore smaller, specialized on-device models to reduce both latency and cost.
@@ A.1. Overview of Computational Load
[…]
@@ A.2. Asynchronous Processing with Multi-Batch Responses
Figure 5. DIY-MOD processing pipeline showing synchronous and asynchronous execution paths. The critical path (diagonal hatching) completes in 5-15 seconds for a batch of posts, returning text modifications and polling tokens. Image transformations process asynchronously in background workers (dotted pattern). The client polls for completed transformations—missing on attempts until the content transformation is ready. Previously processed content can return immediately from cache on first poll. Note that, the critical path delay is for a batch of usually 25 posts in reddit. And, this delay only happens when the user first loads/refreshes the feed in browser. For subsequent batches of posts within the browsing session, predictive prefetching (Section˜A.3) minimizes the delay under reasonable scrolling behavior.
@@
Our polling strategy reflects two key insights. First, analyzing the DOM structure, we found that Reddit employs lazy loading for images, fetching actual image content only when necessary based on viewport position and network conditions. Our deferred image processing aligns naturally with this behavior. Second, the probability that intervention selection has completed for a content increases over time as more candidates are evaluated. Therefore, the client implements an inverse exponential backoff polling strategy, starting with longer intervals that gradually decrease. To prevent server overload from synchronized polling, each content item adds randomized jitter to its polling delay. Posts appearing earlier in the feed receive higher polling priority since users are more likely to encounter them first.
@@ A.4. Content-Based Caching
The system implements a Redis based cache to eliminate redundant processing when multiple users encounter the same content or when users reload pages. The cache key combines three components:
$\mathtt{cache\_key}=\mathtt{hash}(\mathtt{content})\oplus\mathtt{hash}(\mathtt{filter\_params})\oplus\mathtt{hash}(\mathtt{sensitivity\_level})$
This formulation ensures that identical content processed with identical filter parameters yields the same cached result, while different filter configurations or sensitivity levels generate distinct cache entries. The cache operates across all users without compromising privacy since only the anonymous analysis result is stored with no link to user identity.
@@
A.5. Edit Failures, False Obfuscations and Mitigation Strategies
A.5.1. Generation failures and safeguards.
We designed our mitigation strategy to treat both as fail-safe defaults. When generation fails, we first retry, which resolve in most cases. Generating $K=3$ parallel candidates also increases the likelihood that at least one succeeds and scored higher. While we could switch to alternative VLMs upon multiple failure, we prioritized real-time feed rendering over exhaustive fallback strategies. Even when all candidates fail to generate, we leave the content unchanged but still apply the transparency indicator (Section˜4.2, Figure˜1(d)) on top of the original post. This indicator informs users that the content matched one of their filters and was meant to be transformed. Participants reported using this indicator as a personalized warning label—a cue to look away from posts that might be harmful to them. In other words, we err on the side of flagging a post as potentially triggering rather than exposing users to it.
A.5.2. Audit of edit failures and false positives.
To quantify edit failures and false obfuscations, we manually audited posts that our extension marked as modified via the transparency indicator. For each participant, we randomly sampled 5 text posts and 5 image posts, yielding 75 text items and 75 image items. For text, we confirmed via screen recordings that all 75 posts were successfully transformed. For images, we visually matched each transformed image to the original Reddit post and the filter description of the viewer. We found that 66/75 ($\approx 88\%$) images were visually transformed, while 8 ($\approx 12\%$) were either left unchanged or edited in the wrong region. Among the 66 transformed images, 62 ($\approx 93\%$) were correctly transformed according to the viewer’s filter description, implying a false-positive transformation rate of roughly 7% of transformed images.
Appendix B Formative Study
B.1. Data Analysis and Protection
We transcribed the interviews verbatim and then deleted the original recordings to avoid retaining identifiable data. Using reflexive thematic analysis (braun2019reflecting), we inductively identified the core themes presented in Section 3.2 through 3.7. We pseudonymized transcripts with coded identifiers (e.g., P_01, P_12). We stored names and email addresses for recruitment and compensation in a separate encrypted file, deleted them after payment, and never linked them to transcripts. All study data remained on encrypted, university-approved devices and cloud services accessible only to the core research team.
Appendix C More Details About Systems
C.1. DIYMOD
Details in Figure˜5(a) and Figure˜5(b).
@@ C.1. DIYMOD
(5(a)) Chat interface for filter creation.
@@
(5(b)) Filter configuration interface. After establishing the filter description, users configure three parameters: (1) Content Type specifies whether the filter applies to text, images, or both; (2) Sensitivity Level indicates the user’s distress intensity; and (3) Duration sets filter expiration. For brevity, in popup interface we only showed three time presets. Users can modify the filter description anytime using the "Modify Filter" button or add metadata through the Options page for further customization.
C.2. Dual Feed

### v3

@@ 6.2.3. Quantitative Alignment with User Preferences
To complement our qualitative findings, we modeled the probability that participants would select the system’s highest-ranked intervention over a nearby alternative—the randomly selected intervention from ranks two through five. We use an intercept-only model to predict our dependent variable, System Choice Favored, which is a binary variable equal to 1 if the participant selected the system’s choice, and 0 otherwise. We use the glmer function from the lme4 package (Bates et al., 2015) in R. For image interventions, the estimated intercept was positive and statistically significant (log-odds = 0.53, $p=.033$), corresponding to a 63% probability that participants chose the system’s top candidate over the close contenders. For text interventions, the alignment was stronger (log-odds = $1.33$, p<.001{p}<.001), corresponding to a 79.1% probability.
@@ 7.1. Transformation Enables Access, Not Avoidance
Empowering users to take moderation decisions into their own hands might raise concerns about whether this agency can trap them in filter bubbles (Pariser, 2011). DIY-MOD can support a different dynamic. By transforming how content appears rather than removing it entirely, the system enables users to engage with content they would otherwise avoid completely.
@@
At the same time, this approach surfaces deeper tensions: Who decides what users should or should not engage with, and why does that decision need to be made in a top-down, prescriptive manner? The First Amendment is often framed as a protection of speech. Legal scholars have similarly argued for the right to freedom of listening and the right against compelled listening (Corbin, 2009). While these protections do not extend to platforms as private entities, the principles that they represent remain relevant. Our studies show that content that appears benign to most can be experienced as harmful by others. This variability complicates any universal judgment about what content is “safe” or “appropriate”—at least outside the categories codified by law. And it raises a critical question: Who gets to decide which experiences of harm are valid and which users are entitled to protection?
@@ 7.2. The Case for Platform-Integrated Personalization
While DIY-MOD demonstrates personalized moderation is viable as middleware (Fukuyama et al., 2020), its architecture requires intercepting, externally processing, and then scoring generated transformations. These steps introduce unavoidable latency and complicate implementation.
Platforms however, could avoid these bottlenecks entirely. With direct content access, computational resources, and existing user behavioral signals, they could implement personalized moderation far more efficiently. Platforms already use this infrastructure for ad targeting and safety interventions like suicide prevention (Card, 2018; Meta, 2022). The same systems could also power consensual, user-directed personalized content transformation.
@@
Our approach gives rise to a natural question: Which kinds of content should be addressed through platform-led moderation, and which might be better suited to user-driven transformation? Platforms are widely expected to bear responsibility for addressing harms that have broad, collective consequences such as coordinated manipulation campaigns or abuse like hate speech or targeted harassment that fosters hostile environments. Yet even here, the boundaries are far from settled. The same activity can be cast as harassment or as activism, as disinformation or as political dissent. Platforms’ judgments about what to enforce are never neutral. They are shaped by political pressures, business incentives, and cultural norms (Díaz and Hecht-Felella, 2021; Jahanbakhsh, 2023).
@@ 7.4. Application to Civic Discourse
This transformation-based approach may also hold promise for other types of content not brought up by our study participants. One can imagine its application to civic or ideological discourse: when valuable perspectives are obscured by hostile or inflammatory framing, repackaging that content (e.g., by softening antagonistic language, reducing confrontational cues, or highlighting shared values), could make it easier to engage with. It is also worth noting that there is no clear consensus that exposure to counter-attitudinal information is always beneficial. In fact, a body of research shows that such exposure, especially when unfiltered or confrontational, can exacerbate polarization and deepen animosity (Törnberg, 2022; Bail, 2022). This calls into question that more exposure is always better, and highlights the need for systems that support constructive engagement, rather than unmediated confrontation (Bail, 2022). Altering the tone or phrasing of civic or political content through transformations like ours may offer a path toward openness without overwhelm. It is also worth considering who is most likely to adopt such tools. Prior work suggests challenge-averse individuals may use personalization to reinforce existing views (Munson and Resnick, 2010). However, those findings reflect systems that offer a binary choice (e.g., show/hide). By contrast, transformation reduces the affective load of counter-attitudinal content. It is plausible that this, unlike binary filtering, might enable engagement from users who would otherwise disengage completely. Investigating these adoption patterns remains a critical area for future research.
Prior work on bridging divides has largely focused on identifying and ranking content that resonates across ideological lines (Ovadya and Thorburn, 2022; Stray, 2021). While valuable, such approaches depend on the availability of naturally bridgeable posts. But what if we could make more content bridgeable? This of course, raises important questions about authenticity and message integrity: How much transformation is too much? When does reframing a message become misrepresentation? Future work must address these ethical boundaries.
@@ 7.5. Therapeutic Grounding and Applications of Our Work
DIY-MOD’s design is inspired by exposure therapy: enabling safer engagement with a weakened or modified version of a stimulus, rather than avoidance (Rothbaum and Schwartz, 2002; Abramowitz et al., 2019). This principle is reflected in our palette of interventions. Our transformations are designed to function as weakened versions of the original content, and the ability to easily adjust sensitivity levels and filter durations echoes the concept of graduated exposure.
While DIY-MOD is not a clinical intervention in its current form, its combination of weakened content, adjustable sensitivities, and time-bounded filters hints at how similar mechanisms could eventually complement therapeutic practices. For example, these knobs could let users (and clinicians) gradually decrease intervention intensity over time, if they so choose, rather than assuming static sensitivities. This direction aligns with calls (Wang et al., 2025) for practitioner-assisted AI systems that support personalized, adaptive exposure in everyday contexts.
@@ 7.6.1. Architectural Safeguards:
Using AI to interpret personal sensitivities carries inherent risks, as system failures can expose users to content they explicitly sought to avoid. One way we mitigate these risks is by implementing a two-stage cascade architecture: a pruner and then a scorer VLM-judge evaluate content using a consistent rubric. Such multi-stage evaluation strengthens the reliability of LLM/VLM-judge based pipelines (Badshah and Sajjad, 2024; Gu et al., 2024). We manually audited sampled content and found that all sampled text edits and approximately 89% of sampled image edits were successfully transformed, with false obfuscations (edits that did not match the viewer’s filter description) in only about 7% of transformed images (Appendix C).
@@
Privacy considerations are fundamental when handling data tied to personal sensitivities. DIY-MOD adopts a “privacy-by-design” approach (Cavoukian and others, 2009) that prioritizes user control and data minimization. It provides full functionality without an account and offers user-controlled portability of filters through an anonymous export/import feature. By default, it uses local-first storage and practices data minimization, sending only public in-feed content for ephemeral analysis. Looking ahead, portions of the analysis can migrate to lightweight on-device models to further reduce reliance on server-side processing.
@@ 8. Limitations and Future Directions
Improving Selection Model. Our findings reveal that the effectiveness of an intervention is context-dependent, and that successful transformations manage a user’s cognitive state by providing closure. Our system’s selection framework, in contrast, relied on a limited user model derived from filter descriptions and lacked an explicit model of cognitive closure or the user’s immediate context. Improving its ability to select transformations that provide such closure is an important area for future work. The principles and preference data from this study provide a direct path to do so by aligning the selector with human preferences using alignment approaches such as DPO (Rafailov et al., 2023).
Long-Term Psychological Implications. Our study focused on users’ immediate experiences, which raises a longitudinal question: do personalized transformations support long-term wellbeing, or can they enable patterns of unhealthy avoidance? There is good reason to shield users from repeated triggers, as re-exposure may exacerbate anxiety (Markowitz and Fanselow, 2020). At the same time, indefinite avoidance may also have adverse psychological and social effects. Future work should explore how to design for this tension perhaps by incorporating features that allow users to optionally and safely decrease intervention intensity over time, aligning the tool more closely with therapeutic goals.
@@
Technical and Platform Dependencies. While DIY-MOD’s core logic is generalizable, the content interception requires platform-specific adapters, and extending DIY-MOD to video-first (e.g., TikTok) platforms would require additional engineering effort. Further, our reliance on commercial VLMs means their capabilities, biases, and safety filters can affect performance. We detail specific failure modes and mitigations in Appendix C; developing benchmarks to quantify edit failures and over-detection rates remains a potential future work.
Scalability & Latency. Our middleware approach also introduces computational and monetary costs. Our optimizations (detailed in Appendix A) keep the per-batch latency within a usable range for real-time interaction, but high cost associated with LLM/VLM usage remains a scalability constraint. These constraints highlight the efficiency gains of a platform-integrated approach (Section 7.2). Future work could also explore smaller, specialized on-device models to reduce both latency and cost.
@@ 10. Conclusion
Acknowledgements.
We thank Jonathan Stray and Ian Baker for providing access to the codebase for the Prosocial Ranking Challenge. We are grateful to Dr. Elizabeth Duval, Shmeelok Chakraborty, Shwetha Rajaram, Rashidujjaman Rifat, Erika Arias, Rashon Poole, Larnell Moore, and Teanna Sims for their support and feedback throughout the project. Rayhan Rashed also thanks Ehsan Hoque for mentorship that continues to shape how he approaches writing, particularly paper intros.
@@ A.1. Overview of Computational Load
Text Processing Sequence (Critical Path):
 Filter matching NN 0.5-1s Yes
 Intervention selection $M_{t}$ 1-2s Yes
 Intervention application $M_{t}$ 1-2s Yes
Image Filter Analysis (Critical Path, parallel with text):
 Filter matching $N_{p}$ 2-5s Yes
Image Transformation (Asynchronous, Background Task):
 Intervention generation $K\times M_{p}$ 2-10s No
 Intervention scoring $K\times M_{p}$ 2-5s No
@@ A.2. Asynchronous Processing with Multi-Batch Responses
Figure 5. DIY-MOD processing pipeline showing synchronous and asynchronous execution paths. The critical path (diagonal hatching) completes in 5-15 seconds for a batch of posts, returning text modifications and polling tokens. Image transformations process asynchronously in background workers (dotted pattern). The client polls for completed transformations—missing on attempts until the content transformation is ready. Previously processed content can return immediately from cache on first poll. Note that, the critical path delay is for a batch of usually 25 posts in reddit. And, this delay only happens when the user first loads/refreshes the feed in browser. For subsequent batches of posts within the browsing session, predictive prefetching (Section A.3) minimizes the delay under reasonable scrolling behavior.
@@
Our polling strategy reflects two key insights.
First, analyzing the DOM structure, we found that Reddit employs lazy loading for images, fetching actual image content only when necessary based on viewport position and network conditions. Our deferred image processing aligns naturally with this behavior.
Second, the probability that intervention selection has completed for a content increases over time as more candidates are evaluated. Therefore, the client implements an inverse exponential backoff polling strategy, starting with longer intervals that gradually decrease. To prevent server overload from synchronized polling, each content item adds randomized jitter to its polling delay. Posts appearing earlier in the feed receive higher polling priority since users are more likely to encounter them first.
@@ A.4. Content-Based Caching
DIY-MOD implements a Redis based cache to eliminate redundant processing when multiple users encounter the same content or when a page reloads. The cache key combines three components:
$\mathtt{cache\_key}$ $=\mathtt{hash}(\mathtt{content})$
$\quad\oplus\mathtt{hash}(\mathtt{filter\_params})$
$\quad\oplus\mathtt{hash}(\mathtt{sensitivity\_level})$
This formulation ensures that identical content processed with identical filter parameters yields the same cached result. The cache operates across all users without compromising privacy since only the anonymous analysis result is stored with no link to user identity.
@@
(6(a)) Chat interface for filter creation.
@@
(6(b)) Filter configuration interface. After establishing the filter description, users configure three parameters: (1) Content Type specifies whether the filter applies to text, images, or both; (2) Sensitivity Level indicates the user’s distress intensity; and (3) Duration sets filter expiration. For brevity, in popup interface we only showed three time presets. Users can modify the filter description anytime using the ”Modify Filter” button or add metadata through the Options page for further customization.
Appendix B Filter Creation Interface
Details in Figure 6(a) and Figure 6(b).
Appendix C Edit Failures, False Obfuscations and Mitigation Strategies
C.1. Generation failures and safeguards.
We designed our mitigation strategy to treat both as fail-safe defaults. When generation fails, we first retry, which resolve in most cases. Generating $K=3$ parallel candidates also increases the likelihood that at least one succeeds and scored higher. While we could switch to alternative VLMs upon multiple failure, we prioritized real-time feed rendering over exhaustive fallback strategies. Even when all candidates fail to generate, we leave the content unchanged but still apply the transparency indicator (Section 4.2, Figure 1(d)) on top of the original post. This indicator informs users that the content matched one of their filters and was meant to be transformed. Participants reported using this indicator as a personalized warning label—a cue to look away from posts that might be harmful to them. In other words, we err on the side of flagging a post as potentially triggering rather than exposing users to it.
C.2. Audit of edit failures and false positives.
To quantify edit failures and false obfuscations, we manually audited posts that our extension marked as modified via the transparency indicator. For each participant, we randomly sampled 5 text posts and 5 image posts, yielding 75 text items and 75 image items. For text, we confirmed via screen recordings that all 75 posts were successfully transformed. For images, we visually matched each transformed image to the original Reddit post and the filter description of the viewer. We found that 66/75 ($\approx 88\%$) images were visually transformed, while 9 ($\approx 12\%$) were either left unchanged or edited in the wrong region. Among the 66 transformed images, 62 ($\approx 93\%$) were correctly transformed according to the viewer’s filter description, implying a false-positive transformation rate of roughly 7% of transformed images.
Appendix D Formative Study: Data Analysis and Protection
We transcribed the interviews verbatim and then deleted the original recordings to avoid retaining identifiable data. Using reflexive thematic analysis (Braun and Clarke, 2019), we inductively identified the core themes presented in Section 3.2 through 3.7. We pseudonymized transcripts with coded identifiers (e.g., P_01, P_12). We stored names and email addresses for recruitment and compensation in a separate encrypted file, deleted them after payment, and never linked them to transcripts. All study data remained on encrypted, university-approved devices and cloud services accessible only to the core research team.